OpenStudy (lgbasallote):

Two vectors lying in the xy plane are given by the equations \(\vec A = 5\hat i + 2\hat j\) and \(\vec B = 2\hat i - 3\hat j\). The value of \(\vec A \times \vec B\) is what?

OpenStudy (lgbasallote):

i thought it would be \[(5\times 2)\hat i + (2 \times -3)\hat j \implies 10\hat i - 6\hat j\] but it seems im wrong...

OpenStudy (lgbasallote):

@curiousshubham do you know this shiz?

OpenStudy (anonymous):

you are right..... now its 10 - 6 = 4.

OpenStudy (lgbasallote):

well the weird thing is...the choices are \[A. \; 19 \hat k\] \[B \; -11 \hat k\] \[C \; -19 \hat k\] \[D \; 11 \hat k\]

OpenStudy (lgbasallote):

any ideas?

OpenStudy (anonymous):

there's no relation between the unit vector k here....

OpenStudy (lgbasallote):

that's why i say it's weird

OpenStudy (anonymous):

@lgbasallote Where did you find this simple ques with weird options???

OpenStudy (lgbasallote):

same one i have been getting my vector questions from

OpenStudy (lgbasallote):

says the answer is C

OpenStudy (lgbasallote):

no idea @curiousshubham ?

OpenStudy (lgbasallote):

wow

OpenStudy (anonymous):

I was there..... Good Job @demitris ...

OpenStudy (lgbasallote):

so..can you explain what he did?

OpenStudy (lgbasallote):

nevermind..i got it